# Data Mixing Can Induce Phase Transitions in Knowledge Acquisition

### v1

@@ Abstract
Large Language Models (LLMs) are typically trained on data mixtures: most data come from web scrapes, while a small portion is curated from high-quality sources with dense domain-specific knowledge. In this paper, we show that when training LLMs on such data mixtures, knowledge acquisition from knowledge-dense datasets—unlike training exclusively on knowledge-dense data (Allen-Zhu and Li, 2024a)does not always follow a smooth scaling law but can exhibit phase transitions with respect to the mixing ratio and model size. Through controlled experiments on a synthetic biography dataset mixed with web-scraped data, we demonstrate that: (1) as we increase the model size to a critical value, the model suddenly transitions from memorizing very few to most of the biographies; (2) below a critical mixing ratio, the model memorizes almost nothing even with extensive training, but beyond this threshold, it rapidly memorizes more biographies. We attribute these phase transitions to a capacity allocation phenomenon: a model with bounded capacity must act like a knapsack problem solver to minimize the overall test loss, and the optimal allocation across datasets can change discontinuously as the model size or mixing ratio varies. We formalize this intuition in an information-theoretic framework and reveal that these phase transitions are predictable, with the critical mixing ratio following a power-law relationship with the model size. Our findings highlight a concrete case where a good mixing recipe for large models may not be optimal for small models, and vice versa.
@@ 1 Introduction
The pre-training data of large language models (LLMs) can be categorized into two major types. The first type consists of large-scale corpora scraped from the web (Raffel et al., 2020; Penedo et al., 2024; Li et al., 2024), often spanning billions to trillions of tokens across diverse topics and styles. Due to the scale, it is inherently hard to ensure the information density of the dataset and its relevance to downstream tasks. Hence, a second type of data, smaller-scale datasets curated from high-quality sources, is incorporated. This type of data usually contains very dense knowledge on tasks or domains with significant practical value. For example, Wikipedia and Stack Exchange cover a wide range of world knowledge. OpenWebMath (Paster et al., 2024) and StarCoder (Li et al., 2023; Kocetkov et al., 2022) provide valuable data for improving model performance on mathematics and coding tasks.
@@
The second type of data, which we refer to as knowledge-dense data, typically accounts for only a small fraction of the entire corpus. In the pre-training data of a recently released model family, OLMo 2 (OLMo et al., 2025), over 95% of the tokens are from web data, and only less than 5% are from knowledge-dense data. The proportion of each individual knowledge-dense dataset is even smaller, e.g., only less than 0.1% of the tokens are from Wikipedia. This naturally raises a question: How much knowledge can LLMs really acquire from this small amount of knowledge-dense data?
@@
If LLMs were exclusively trained on knowledge-dense data without any data mixing, the amount of knowledge acquired after sufficient training should scale linearly with model size. Although quantifying knowledge in natural data is non-trivial, Allen-Zhu and Li (2024a) sidestep this issue and provide strong empirical evidence for this linear scaling law through extensive pre-training experiments on synthetically generated biographies. In their setting, the amount of knowledge stored by a model is quantified by evaluating its memorization of the biographies using information-theoretic metrics. Similar linear scaling laws are also observed in memorizing Wikidata fact triples by Lu et al. (2024), and analyzed theoretically by Nichani et al. (2025). Based on these results, one might naively expect a similar linear relationship between model size and acquired knowledge when knowledge-dense data is mixed with web data.
@@
However, in this paper, we show that the linear scaling no longer holds under data mixing. We consider the setup where a knowledge-dense dataset focused on a single domain constitutes a small fraction $r$ of the pre-training corpus—referred to as the mixing ratio—and the rest is large-scale web text (see Section C.1 for our implementation of data mixing). We demonstrate via a quantitative study that knowledge acquisition from the knowledge-dense data exhibits a more intricate behavior with notable phase transitions with respect to the mixing ratio and model size.
@@
More specifically, we study factual knowledge acquisition. We follow the approach of Allen-Zhu and Li (2024a) to curate a synthetic dataset of biographies, where each individual’s information is embedded into natural text descriptions using diverse templates. Due to the uniform data format and content of this dataset, we can quantify how much knowledge the model has stored simply by counting the number of memorized biographies. We then mix this synthetic biography dataset with large-scale web corpus FineWeb-Edu (Penedo et al., 2024) or the Pile (Gao et al., 2020) to create the pre-training mixture. We pre-train or continually pre-train Pythia models (Biderman et al., 2023) ranging from 14M to 6.9B parameters on these mixtures.
@@
Finding 2: Phase Transition in Mixing Ratio (Figures 2 and 10).
@@
    Naively measuring the performance of small models on a small data domain may provide little to no predictive signal on how well larger models perform, revealing a potential limitation of using small proxy models for data curation, as also evidenced by Kang et al. (2024); Jiang et al. (2024); Ye et al. (2024); Magnusson et al. (2025).
@@ 2 Experimental Setup
We follow  Allen-Zhu and Li (2024b) to create a synthetic biography dataset, with each individual characterized by five attributes: birth date, birth city, university, major, and employer. For each individual, the value of each attribute is randomly and independently sampled from a predefined domain. These (name, attribute, value) triplets are then converted into natural text using sentence templates. For instance, (Gracie Tessa Howell, birth city, St. Louis, MO) is converted into “Gracie Tessa Howell’s birthplace is St. Louis, MO.” Following (Allen-Zhu and Li, 2024b), every time the model encounters a biography, the five sentences are randomly shuffled, and a new sentence template is selected for each attribute from a set of five possible templates. We denote the dataset containing $N$ biographies as SynBio-$N$. See Section C.2.1 for full details.
@@
Denote a knowledge triplet (name, attribute, value) as $(\bm{n},\bm{a},\bm{v})$ and let $\lvert\bm{v}\rvert$ represent the number of tokens in $\bm{v}$. For evaluation, the model is prompted with the sentence prefix containing $\bm{n}$ and $\bm{a}$ and is tasked to generate $\lvert\bm{v}\rvert$ tokens via greedy decoding. A triplet is considered learned if the output exactly matches $\bm{v}$. For example, given the triplet (Gracie Tessa Howell, birth city, St. Louis, MO), the prompt “Gracie Tessa Howell’s birthplace is” is provided. We say the model has learned the fact if it generates “St. Louis, MO.” We report the accuracy averaged over all individuals, attributes, and templates in the main text and defer the detailed results to Section B.4.
@@
Our experiments use the Pythia architecture (Biderman et al., 2023), with model sizes ranging from 14M to 6.9B. The default setup involves pre-training from scratch on a mixture of FineWeb-Edu and SynBio. Since FineWeb-Edu is large ($>$1T tokens) and SynBio is small ($<$1B tokens), our typical training runs involve the model seeing SynBio for multiple epochs but FineWeb-Edu for less than one epoch. For instance, in a 32B-token run with the mixing ratio for SynBio-320k set as 0.1, the model passes SynBio $\sim 100$ times. We also study the continual pre-training setup in Sections 6 and B.1. Full details are provided in Appendix C.
@@ 3.1 Phase Transition in Model Size
We first investigate how knowledge acquisition within data mixtures is affected by model size at fixed mixing ratios. For each $r\in\{0.1,0.2,0.3,0.4\}$, we train models with sizes from 14M to 410M on the mixture of FineWeb-Edu and SynBio-320k for a sufficiently long horizon of 32B tokens, which is approximately four times the optimal computation for 410M models predicted by the Chinchilla scaling law (Hoffmann et al., 2022). As shown in Figure 2, as model size increases, accuracy on SynBio initially remains near zero. Once the model size surpasses some threshold, accuracy rapidly grows to above 60%. The transition is consistently sharp across different mixing ratios while larger $r$ leads to a smaller critical point.
@@ 3.2 Phase Transition in Mixing Ratio
We begin by training models of the same size with different mixing ratios $r$. Specifically, we train 70M models on the mixture of FineWeb-Edu and SynBio-320K, varying $r$ from 0.1 to 0.45 (stepsize 0.05), and 410M models on the mixture of FineWeb-Edu and SynBio-1.28M, varying $r$ from 0.1 to 0.4 (stepsize 0.1). All models are trained for a total of 32B tokens. As shown in  Figure 2(a), for 70M models, as $r$ increases from 0.1 to 0.25, its accuracy on SynBio remains near zero. Only when $r>0.3$ does the accuracy begin to steadily improve. In Figure 2(b), the accuracy for 410M models exhibit similar trends where it remains near zero for $r\leq 0.3$ and suddenly attains 80% when $r$ grows to 0.4. In Figure 10, we replicate the experiments on Pythia 2.8B and 6.9B models to show that similar phase transition in mixing ratio persists for larger models.
@@
Given the observed phase transition, one may raise the following counter-argument: if models are trained for a sufficiently long horizon—such that even a small mixing ratio $r$ would eventually result in each biography being encountered hundreds or even thousands of times—then the phase transition might no longer exist. To test this counter-argument, we extend the training horizon for $r=0.2$ to 512B tokens for the 70M and 410M models by 16 and 4 times respectively. Under this extended training, each biography appears $\sim 200$ times for the 70M model and $\sim 3000$ times for the 410M model. As shown in Figures 3(a) and 6, the accuracy on SynBio remains near zero even after such extensions.
@@
3.3 Ablation Studies
(a) Vary the batch size.
(b) Vary the peak learning rate.
Consistent Trends Across Different Batch Sizes.
Consistent trends across learning rate values and schedules.
@@ 3.4 Phase Transitions on Reasoning Tasks
In this subsection, we show that phase transitions can also arise when mixing a knowledge-dense dataset aimed at improving models’ reasoning skills with web text. Such datasets are often multi-task in practice. For example, OpenWebMath (Paster et al., 2024) covers diverse math topics. We show that phase transitions can occur for each single subtask within this knowledge-dense dataset. Inspired by Ruis et al. (2024), we consider slope calculation between two points $(x_{1},y_{1})$ and $(x_{2},y_{2})$, and replace all the documents containing the word “slope” in OpenWebMath with our cleaner and higher quality slope data. We then mix the modified OpenWebMath with FineWeb-Edu and train Pythia models from scratch. Similar to the setup of SynBio, every time the model sees a slope calculation example, we uniformly sample $x_{1},y_{1},x_{2},y_{2}$ from $\{0,1,\cdots,99\}$ (ensuring $x_{1}\neq x_{2}$), and apply randomly chosen question and step-by-step answer templates. For evaluation, we randomly generate 1k questions for slope calculation and check if the model produces the correct final answer. Results in in Figure 6 show similar phase transitions as factual knowledge acquisition. See details in Section C.3.
@@ 4.2 Problem Formulation
¯$\bar{\mathcal{L}}_{\mathcal{P}}(\mathcal{A}):=\mathbb{E}_{\theta\sim\mathcal{P}}[\mathcal{L}(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta})].$ (1)
@@
We measure a model’s “effective” capacity—the amount of information a model, produced by some learning algorithm $\mathcal{A}$, stores about the data distribution $\mathcal{D}_{\theta}$—by the mutual information (MI) between the model and the data distribution $\mathcal{D}_{\theta}$, i.e., $I(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta})$. For practical learning algorithms with bounded capacity, if $\mathcal{A}$ always outputs a model $h$ with at most $N$ parameters each represented by a $b$-bit floating number, then $I(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta})\leq bN$ by information theory. Empirically, Allen-Zhu and Li (2024a) found that $I(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta})\approx 2N$ holds across various training setups by controlled experiments.
@@
where the infimum is taken over all learning algorithms. An optimal $M$-bounded-capacity learner is a learning algorithm $\mathcal{A}$ such that $I(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta})\leq M$ and ¯$\bar{\mathcal{L}}_{\mathcal{P}}(\mathcal{A})=F_{\mathcal{P}}(M)$.
@@ 4.3 Warmup: Training Exclusively on Mixture of Facts
[…]

### v2

@@ Abstract
Large Language Models (LLMs) are typically trained on data mixtures: most data come from web scrapes, while a small portion is curated from high-quality sources with dense domain-specific knowledge. In this paper, we show that when training LLMs on such data mixtures, knowledge acquisition from knowledge-dense datasets—unlike training exclusively on knowledge-dense data (allen2024physics)does not always follow a smooth scaling law but can exhibit phase transitions with respect to the mixing ratio and model size. Through controlled experiments on a synthetic biography dataset mixed with web-scraped data, we demonstrate that: (1) as we increase the model size to a critical value, the model suddenly transitions from memorizing very few to most of the biographies; (2) below a critical mixing ratio, the model memorizes almost nothing even with extensive training, but beyond this threshold, it rapidly memorizes more biographies. We attribute these phase transitions to a capacity allocation phenomenon: a model with bounded capacity must act like a knapsack problem solver to minimize the overall test loss, and the optimal allocation across datasets can change discontinuously as the model size or mixing ratio varies. We formalize this intuition in an information-theoretic framework and reveal that these phase transitions are predictable, with the critical mixing ratio following a power-law relationship with the model size. Our findings highlight a concrete case where a good mixing recipe for large models may not be optimal for small models, and vice versa.
@@ 1 Introduction
The pre-training data of large language models (LLMs) can be categorized into two major types. The first type consists of large-scale corpora scraped from the web (raffel2020c4; penedo2024fineweb; li2024datacomp), often spanning billions to trillions of tokens across diverse topics and styles. Due to the scale, it is inherently hard to ensure the information density of the dataset and its relevance to downstream tasks. Hence, a second type of data, smaller-scale datasets curated from high-quality sources, is incorporated. This type of data usually contains very dense knowledge on tasks or domains with significant practical value. For example, Wikipedia and Stack Exchange cover a wide range of world knowledge. OpenWebMath (paster2024openwebmath) and StarCoder (li2023starcoder; kocetkov2022stack) provide valuable data for improving model performance on mathematics and coding tasks.
@@
The second type of data, which we refer to as knowledge-dense data, typically accounts for only a small fraction of the entire corpus. In the pre-training data of a recently released model family, OLMo 2 (olmo20252olmo2furious), over 95% of the tokens are from web data, and only less than 5% are from knowledge-dense data. The proportion of each individual knowledge-dense dataset is even smaller, e.g., only less than 0.1% of the tokens are from Wikipedia. This naturally raises a question: How much knowledge can LLMs really acquire from this small amount of knowledge-dense data?
@@
If LLMs were exclusively trained on knowledge-dense data without any data mixing, the amount of knowledge acquired after sufficient training should scale linearly with model size. Although quantifying knowledge in natural data is non-trivial, allen2024physics sidestep this issue and provide strong empirical evidence for this linear scaling law through extensive pre-training experiments on synthetically generated biographies. In their setting, the amount of knowledge stored by a model is quantified by evaluating its memorization of the biographies using information-theoretic metrics. Similar linear scaling laws are also observed in memorizing Wikidata fact triples by lu2024scaling, and analyzed theoretically by nichani2025understanding. Based on these results, one might naively expect a similar linear relationship between model size and acquired knowledge when knowledge-dense data is mixed with web data.
@@
However, in this paper, we show that the linear scaling no longer holds under data mixing. We consider the setup where a knowledge-dense dataset focused on a single domain constitutes a small fraction $r$ of the pre-training corpus—referred to as the mixing ratio—and the rest is large-scale web text (see Section E.1 for our implementation of data mixing). We demonstrate via a quantitative study that knowledge acquisition from the knowledge-dense data exhibits a more intricate behavior with notable phase transitions with respect to the mixing ratio and model size.
@@
More specifically, we study factual knowledge acquisition. We follow the approach of allen2024physics to curate a synthetic dataset of biographies, where each individual’s information is embedded into natural text descriptions using diverse templates. Due to the uniform data format and content of this dataset, we can quantify how much knowledge the model has stored simply by counting the number of memorized biographies. We then mix this synthetic biography dataset with large-scale web corpus FineWeb-Edu (penedo2024fineweb) or the Pile (gao2020pile) to create the pre-training mixture. We pre-train or continually pre-train Pythia models (biderman2023pythia) ranging from 14M to 6.9B parameters on these mixtures.
@@
Finding 2: Phase Transition in Mixing Ratio (Figures 2 and 9).
@@
In Figure 10, we further show that the observed phase transitions are not limited to discrete metrics like accuracy, but also persist in validation loss, a continuous metric.
@@
    Naively measuring the performance of small models on a small data domain may provide little to no predictive signal on how well larger models perform, revealing a potential limitation of using small proxy models for data curation, as also evidenced by kang2024autoscale; jiang2024adaptive; ye2024datamixinglaws; magnusson2025datadecide; mizrahi2025languagebetr.
@@ 2 Experimental Setup
We follow  allenzhu2024physicslanguagemodels32 to create a synthetic biography dataset, with each individual characterized by five attributes: birth date, birth city, university, major, and employer. For each individual, the value of each attribute is randomly and independently sampled from a predefined domain. These (name, attribute, value) triplets are then converted into natural text using sentence templates. For instance, (Gracie Tessa Howell, birth city, St. Louis, MO) is converted into “Gracie Tessa Howell’s birthplace is St. Louis, MO.” Following (allenzhu2024physicslanguagemodels32), every time the model encounters a biography, the five sentences are randomly shuffled, and a new sentence template is selected for each attribute from a set of five possible templates. We denote the dataset containing $N$ biographies as SynBio-$N$. See Section E.2.1 for full details.
@@
Denote a knowledge triplet (name, attribute, value) as $(\bm{n},\bm{a},\bm{v})$ and let $\lvert\bm{v}\rvert$ represent the number of tokens in $\bm{v}$. For evaluation, the model is prompted with the sentence prefix containing $\bm{n}$ and $\bm{a}$ and is tasked to generate $\lvert\bm{v}\rvert$ tokens via greedy decoding. We then check whether the output exactly matches $\bm{v}$. For example, given the triplet (Gracie Tessa Howell, birth city, St. Louis, MO), the prompt “Gracie Tessa Howell’s birthplace is” is provided. We say the model has memorized the fact if it generates “St. Louis, MO.” We report the accuracy averaged over all individuals, attributes, and templates in the main text and defer the detailed results to Section D.6.
@@
Our experiments use the Pythia architecture (biderman2023pythia), with model sizes ranging from 14M to 6.9B. The default setup involves pre-training from scratch on a mixture of FineWeb-Edu and SynBio. Since FineWeb-Edu is large (>>1T tokens) and SynBio is small (<<1B tokens), our typical training runs involve the model seeing SynBio for multiple epochs but FineWeb-Edu for less than one epoch. This approach reflects real-world practices to rephrase knowledge-dense data multiple times (hao2025reformulation; team2025kimi), since models struggle to acquire knowledge within a single pass (huang2024demystifying). For instance, in a 32B-token run with the mixing ratio for SynBio-320k set as 0.1, the model passes SynBio $\sim 100$ times. We also study the continual pre-training setup in Sections 6 and D.1. Full experimental details are provided in Appendix E.
@@ 3.1 Phase Transition in Model Size
We first investigate how knowledge acquisition is affected by model size given the data mixture. For each $r\in\{0.1,0.2,0.3,0.4\}$, we train models with sizes from 14M to 410M on the mixture of FineWeb-Edu and SynBio-320k for a sufficiently long horizon of 32B tokens, which is approximately four times the compute-optimal training tokens for 410M models according to hoffmann2022an. As shown in Figure 2, as the model size increases, accuracy on SynBio initially remains near zero. Once the model size surpasses some threshold, accuracy rapidly grows to above 60%. The transition is consistently sharp across different mixing ratios while larger $r$ leads to a smaller critical point.
@@ 3.2 Phase Transition in Mixing Ratio
We begin by training models of the same size with different mixing ratios $r$. Specifically, we train 70M models on the mixture of FineWeb-Edu and SynBio-320K, varying $r$ from 0.1 to 0.45 (stepsize 0.05), and 410M models on the mixture of FineWeb-Edu and SynBio-1.28M, varying $r$ from 0.1 to 0.4 (stepsize 0.1). All models are trained for a total of 32B tokens. As shown in  Figure 2(a), for 70M models, as $r$ increases from 0.1 to 0.25, its accuracy on SynBio remains near zero. Only when $r>0.3$ does the accuracy begin to steadily improve. In Figure 2(b), the accuracy for 410M models exhibit similar trends where it remains near zero for $r\leq 0.3$ and suddenly attains 80% when $r$ grows to 0.4. In Figure 9, we replicate the experiments on Pythia 2.8B and 6.9B models to show that similar phase transition in mixing ratio persists for larger models. In Table 1, we report the mean and standard deviation of accuracy for experiments in Figure 2(a).
@@
Given the observed phase transition, one may raise the following counter-argument: if models are trained for a sufficiently long horizon—such that even a small mixing ratio $r$ would eventually result in each biography being encountered hundreds or even thousands of times—then the phase transition might no longer exist. To test this counter-argument, we extend the training horizon for $r=0.2$ to 512B tokens for the 70M and 410M models by 16 and 4 times respectively. Under this extended training, each biography appears $\sim 3000$ times for the 70M model and $\sim 200$ times for the 410M model. As shown in Figures 3(a) and 5, the accuracy on SynBio remains near zero even after such extensions.
@@
We also conduct ablation studies on hyperparameters in Section D.2.
@@ 3.3 Phase Transitions on Reasoning Tasks
In this subsection, we show that the phase transition phenomenon is not limited to factual knowledge, but also extends to datasets related to reasoning. Such datasets are often multi-task in practice. For example, OpenWebMath (paster2024openwebmath) covers diverse math topics. We show that phase transitions can occur for each single subtask within this dataset. Inspired by ruis2024procedural, we consider the slope calculation task between two points $(x_{1},y_{1})$ and $(x_{2},y_{2})$. To explicitly control the frequency and format of the slope calculation examples, we replace all the documents containing the word “slope” in OpenWebMath with our clean and high-quality slope calculation demonstrations. We then mix the modified OpenWebMath with FineWeb-Edu and train Pythia models on this mixture from scratch. Similar to the setup of SynBio, every time the model sees a slope calculation example, we uniformly sample $x_{1},y_{1},x_{2},y_{2}$ from $\{0,1,\cdots,99\}$ (ensuring $x_{1}\neq x_{2}$), and apply randomly chosen question-answer templates. For evaluation, we randomly generate 1k questions for slope calculation and check if the model produces the correct final answer. Results in Figure 5 show similar phase transitions as factual knowledge acquisition (see details in LABEL:sec:slope-detail). Section D.3 presents further discussions and experiments on another reasoning task with larger input space.
@@ 4.2 Problem Formulation
$\bar{\mathcal{L}}_{\mathcal{P}}(\mathcal{A}):=\mathbb{E}_{\theta\sim\mathcal{P}}[\mathcal{L}(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta})].$ (1)
@@
We measure a model’s “effective” capacity—the amount of information a model, produced by some learning algorithm $\mathcal{A}$, stores about the data distribution $\mathcal{D}_{\theta}$—by the mutual information (MI) between the model and the data distribution $\mathcal{D}_{\theta}$, i.e., $I(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta})$. For practical learning algorithms with bounded capacity, if $\mathcal{A}$ always outputs a model $h$ with at most $N$ parameters each represented by a $b$-bit floating number, then $I(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta})\leq bN$ by information theory. Empirically, allen2024physics found that $I(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta})\approx 2N$ holds across various training setups by controlled experiments.
@@
where the infimum is taken over all learning algorithms. An optimal $M$-bounded-capacity learner is a learning algorithm $\mathcal{A}$ such that $I(\mathcal{A}(\mathcal{D}_{\theta});\mathcal{D}_{\theta})\leq M$ and $\bar{\mathcal{L}}_{\mathcal{P}}(\mathcal{A})=F_{\mathcal{P}}(M)$.
@@ 4.3 Warmup: Training Exclusively on Mixture of Facts
We start with a simple case where the data distribution $\mathcal{D}_{\theta}$ contains $K$ random facts. Each fact is a pair $(X_{i},y_{i})$, where $X_{i}$ is a set of input contexts (e.g., paraphrases) and $y_{i}$ is the target token. For instance, the fact “Gracie Tessa Howell was born in 1946” can have contexts like “Gracie Tessa Howell’s birth year is” or “Gracie Tessa Howell came to this world in the year,” all mapping to the target $y=\text{``1946''}$. We further assume that $X_{1},\dots,X_{K}$ are disjoint.